OpenStudy (anonymous):

can anyone please help me i couldn't figure it out i really need help on this one!!!!!!!!!! i have been trying ti figure it out but it won't work out for me.

OpenStudy (anonymous):

this is a pic of the problem

OpenStudy (anonymous):

i am really good at doing this but i can't figure it out

OpenStudy (anonymous):

would it be $200,000>.6>1/12 arrow is dot in middle

OpenStudy (jdoe0001):

I'd think so, yes

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

would the next part be 120,000>1/12 arrow is actually a dot in the middle

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

so what would like be the answer this is what i have problems on

OpenStudy (jdoe0001):

\(\bf i=prt\implies i=200,000\cdot 0.6\cdot \cfrac{1}{12}\implies i=120,000\cdot \cfrac{1}{12}\)

OpenStudy (anonymous):

in this there is 3 parts i just answered 2 of them but the 3rd i need help on

OpenStudy (anonymous):

which is the answer

OpenStudy (jdoe0001):

hmm the fraction? well.. \(\bf i=prt\implies i=200,000\cdot 0.6\cdot \cfrac{1}{12}\implies i=120,000\cdot \cfrac{1}{12}\\ \quad \\ i=120,000\cdot \cfrac{1}{12}\implies i=\cfrac{120,000}{12}\implies i=10,000\)

OpenStudy (anonymous):

oh that makes scence now thank you

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

gonna ask another

OpenStudy (jdoe0001):

ok